### v1

@@ 3.3.1 M-Encoder (Multi Scale Attention Encoder)
3.3.2 MSD-Bottleneck (Multi-Scale Dilated Bottleneck)
Continuous convolution and pooling operations frequently degrade fine-grained details essential for precise segmentation, while conventional up-sampling methods fail to adequately recover these features. To overcome these limitations, we introduce the MSD-Bottleneck module (Fig. 4), building upon [45]. This module integrates a Bottleneck structure with an Atrous Spatial Pyramid Pooling (ASPP) mechanism to simultaneously preserve fine details and capture multi-scale context. The MSD-Bottleneck employs three progressive dilation patterns ([1,2,1], [2,4,2], [4,8,4]) to efficiently extract multi-scale features without substantially increasing parameters. Residual connections mitigate the vanishing gradient problem, facilitating deeper network training. The ASPP component further enhances contextual understanding through parallel dilated convolutions (rates= [4,8]) and global average pooling. These features are concatenated and projected onto the original channel dimension, effectively combining local precision with global context. Mathematically, for the input $x$, the module first processes it through three dilated bottleneck blocks with residual connections, as formalized in Eq. (2). The output then passes through the ASPP stage.
@@ 3.3.2 MSD-Bottleneck (Multi-Scale Dilated Bottleneck)
$\text{out}_{2}=\text{conv}_{3\times 3}(x;\text{dilation}=[2,4,2]),$
$\text{out}_{3}=\text{conv}_{3\times 3}(x;\text{dilation}=[4,8,4]),$
$\text{out}_{4}=\text{Cat}\Big{(}\text{conv}_{3\times 3}(\text{out}_{3};\text{dilation}=[4,8]),\text{AAP}(\text{out}_{3})\Big{)},$
@@
where, $x$ represents the input feature maps to the MSD-Bottleneck,the notation dilation=[1, 2, 1] specifies three convolutional layers with dilation rates of 1, 2, and 1, respectively, ASPP represents the output feature maps, Cat denotes the concatenation operation in the channel dimension, and AAP represents the Adaptive Average Pooling.
@@ 3.3.3 Contextual Attention Fusion Module (CAFM)
Encoders in segmentation networks hierarchically extract feature maps through progressive downsampling; however, deeper layers often lose critical spatial context, parti- cularly in complex medical imaging such as coronary digital subtraction angiography (DSA). To address this, the Contextual Attention Fusion Module (CAFM) (Fig. 5) enhances feature maps by integrating multi-scale contextual information using dilated convolutions. These convolutions expand the receptive field while preserving resolution, enabling robust feature extraction across varying vessel sizes; a key requirement in coronary DSA. Furthermore, CAFM incorporates Squeeze-and-Excitation (SE) blocks to dynamically recalibrate channel-wise feature importance, ensuring the model prioritizes diagnostically relevant features. By refining feature representations, CAFM supplies the decoder with high-quality, context-rich inputs for precise segmentation reconstruction. Additionally, it bridges the encoder-decoder hierarchy semantic gaps between layers through enhanced multi-scale representations. The module’s ability to capture both fine and coarse anatomical structures significantly improves segmentation performance in medical imaging tasks. Its mathematical formulation is presented in Eq. (3). Let $x$ be the input feature map to the CAFM module. The operations performed by the module can be described as follows, with dilation rates of [1, 2, 4, 8]:
$\text{out}_{\text{final}}=\text{conv}_{3\times 3}\Big{(}SE\big{(}\text{x};\text{dilation}=[1,2,4,8]\big{)}\Big{)}$ (3)
@@
Figure 5: Architecture of the Contextual Attention Fusion Module (CAFM) in the MSA-UNet3+ model, combining Squeeze-and-Excitation (SE) blocks with multi-scale processing.
@@ 3.3.4 Loss function
In this study, we optimize our deep learning model for coronary DSA image segmentation using a composite loss function. Given the challenges of class imbalance and high intra-class variability inherent in such tasks, the choice of loss functions is critical. Although Binary Cross-Entropy (BCE) and Dice loss are commonly employed, they do not inherently enforce semantically meaningful or discriminative feature learning in the encoder. To address this limitation and improve segmentation performance, we propose a novel Supervised Prototypical Contrastive Loss (SPCL), used alongside with BCE and Dice loss. By operating on embedded feature vectors, SPCL encourages the encoder to generate semantically rich and discriminative representations, thereby enhancing the overall performance of the segmentation network.
@@
    is a widely used loss function for binary segmentation tasks. It measures the dissimilarity between the predicted probability map and the ground truth labels. For a binary segmentation task, the BCE loss is defined as in Eq.  (4):
@@
    where $y_{i}$ is the ground truth label (0 or 1) for pixel $i$, $\hat{y}_{i}$ is the predicted probability of pixel $i$ belonging to the foreground class, and $N$ is the total number of pixels. BCE loss ensures the model learns to classify each pixel correctly by penalizing deviations from the ground truth. However, it may struggle with class imbalance, as it treats all pixels equally, regardless of class distribution.
@@
    is a commonly used metric for evaluating segmentation performance, particularly in medical imaging tasks where precise overlap between predicted and ground truth masks is critical. It measures the similarity between the predicted segmentation mask and the ground truth mask. By maximizing this overlap, Dice loss encourages the model to focus on regions of interest, thereby improving segmentation performance. The Dice loss is defined as in Eq.  (5):
@@
    where $y_{i}$ and $\hat{y}_{i}$ are the ground truth and predicted probability for pixel $i$, respectively. Dice loss complements BCE by emphasizing correct foreground class predictions, crucial for precise segmentation in coronary DSA images.
@@
    To address class imbalance and high intra-class variance, we introduce the Supervised Prototypical Contrastive Loss (SPCL). This loss combines supervised and prototypical contrastive losses, enhancing the encoder’s ability to differentiate between diverse classes.
    (1) Supervised Contrastive Loss: This loss encourages the encoder to map feature vectors of similar image samples (foreground or background regions) closer in the embedding space while pushing dissimilar ones apart. The loss function is defined as in Eq. (6):
@@
    where $f_{i}$ and $f_{j}$ are feature vectors, $P(i)$ is the set of positive pairs for $i$, $A(i)$ is the set of all pairs for $i$, sim is a cosine similarity function, and $\tau$ is a temperature parameter set to 1 without hyperparameter tuning in this work.
@@
    where, $N$: Total number of valid embeddings (pixels), $n_{\text{p}}$: Number of prototypes, $D\,(z_{i},p_{k})$: Cosine distance between embedding $z_{i}$ and prototype $p_{k}$, $w_{1}$: Weight for positive samples (close to their prototype), $w_{0}$: Weight for negative samples (far from their prototype), and $m$: Margin parameter controlling class separation.
@@ 4.1 Dataset
The DSA dataset used in this study was obtained from the authors of [46], with original data sourced from the General Hospital of Southern Theatre Command. It comprises 300 coronary X-ray angiographic sequences from 50 patients, each with a resolution of $512\times 512$ pixels, including left and right coronary imaging sessions. This study uses only the right coronary images (150 images). The videos capture the flow of contrast agent through the coronary artery, and clinicians select the most clinically significant frames. After pre-processing, images were resized to $256\times 256$ pixels. The dataset is divided into 120 training images and 30 testing images. For training, 5-fold cross-validation is applied, dividing the 120 training images into five equal folds. The results are averaged across all folds for the final evaluation.
@@ 4.2 Implementation Details
The proposed model was implemented using PyTorch and PyTorch Lightning, providing a flexible and scalable en- vironment for deep learning experiments. The experiments were carried out on an NVIDIA RTX 4090 GPU with 24 GB of memory. The dataset was organized into training and test sets, with images and labels resized to $256\times 256$ pixels. Pre-processing included normalization to the range $[-1,1]$ using a mean and standard deviation of 0.5 [47]. Data augmentation techniques, such as random brightness [48] and contrast adjustments [49], were applied during training to enhance generalization. The model was trained using the Adam optimizer with an initial learning rate of 0.0001 and a multi-step scheduler reducing the rate by a factor of 0.1 at 60% and 80% of the training epochs. Training spanned 100 epochs with a batch size of 5.
@@ 4.3 Evaluation Metrics
    It measures the overlap between the predicted segmentation mask and the ground truth. Widely used in medical image segmentation, it emphasizes correct foreground class predictions. The mathematical representation of DSC is illustrated in Eq. ( 9).
@@
    where, True Positives (TP): Pixels correctly predicted as foreground; False Positives (FP): Pixels incorrectly predicted as foreground; and False Negatives (FN): Pixels incorrectly predicted as background.
@@
    Table 1: Comparative results of segmentation models on the coronary DSA dataset, evaluating the effectiveness of the proposed SPCL. Performance metrics for six architectures are presented for ($\gamma=0$ (baseline) and $\gamma=1$), with means and standard deviations from 5-fold cross-validation.
@@
  3. 3.
    F1-Score
[…]
    Precision measures the proportion of correctly predicted foreground pixels ($\text{Precision}=\frac{TP}{FP+TP}$). The F1-score balances precision and recall, making it a comprehensive metric for evaluating segmentation performance, particularly in imbalanced datasets.
  4. 4.
    Average Surface Distance (ASD)
    ASD measures the average distance between the surfaces of the predicted and ground-truth masks, as defined in Eq.  (12):
    where, $S_{P}$ and $S_{G}$ are the surfaces of the predicted and ground-truth masks, respectively; $d(p,S_{G})$ is the minimum Euclidean distance from point $p$ on $S_{P}$ to $S_{G}$; and $d(g,S_{P})$ is the minimum Euclidean distance from point $g$ on $S_{G}$ to $S_{P}$.
    ASD evaluates segmentation boundary accuracy, with lower values indicating better alignment between predicted and ground-truth contours, crucial for tasks such as coronary artery segmentation.
  5. 5.
    Average Contour Distance (ACD)
    ACD measures the average distance between the contours of the predicted and ground-truth masks, specifically focusing on contour points. It is defined as in Eq. (13):
@@ 4.4 Comparison Results
As shown in Table 1, our 5-fold cross-validated results on the coronary DSA dataset demonstrate consistent and statistically significant improvements (one-tailed t-test, p < 0.05) when employing SPCL. Specifically, the Dice coefficient increases from 87.33%, 87.49%, 86.81%, 86.21%, 87.15%, and 85.68% to 87.87%, 87.85%, 87.83%, 87.03%, 87.79%, and 86.80% for U-Net, Att U-Net, UNet++, R2U-Net, Attention UNet++, and R2AU-Net respectively. These quantitative improvements across all architectures and metrics confirm SPCL’s effectiveness as a generalizable enhancement for medical image segmentation tasks.
Figure 7: Visualization comparison of R2AU-Net segmentation performance with and without SPCL loss on coronary DSA images. From left to right: Original angiogram, Ground truth annotation, Baseline segmentation ($\gamma=0$), and SPCL-enhanced results ($\gamma=1$). Yellow rectangles highlight regions where SPCL improves vessel continuity and reduces false positives, particularly in complex branching areas and low-contrast regions.
Table 2: Performance comparison of state-of-the-art medical image segmentation models on the coronary DSA dataset. Results are presented as means $\pm$ standard deviations across 5-fold cross-validation. Statistical significance of performance improvements achieved by MSA-UNet3+ is verified using a one-tailed $t$-test.
@@
Our comparative analysis of loss functions for coronary DSA segmentation (Fig. 6) reveals that integrating SPCL with Dice+BCE yields consistent improvements in all architectures, with DSC gains ranging from +0.36 (U-Net) to +1.12 (R2AU-Net). This performance progression aligns with the findings presented in Table 1, confirming the dual ability of SPCL to: (1) improve feature discrimination through semantic clustering and (2) address class imbalance through contrastive learning. The demonstrated robustness across architectures (U-Net to R2AU-Net) and the magnitude of improvement position SPCL as a general solution for medical segmentation tasks.
These results corroborate our earlier quantitative assessments while providing new insights into SPCL’s architecture-dependent optimization characteristics. The qualitative comparison in Fig. 7 demonstrates SPCL’s effectiveness through two key improvements: (1) enhanced vessel continuity, particularly in complex branching structures where the baseline model fragments vessels; and (2) superior preservation of thin vessels that conventional approaches often miss.
These visual improvements align with our quantitative metrics as illustrated in Table 1. The yellow-highlighted regions specifically showcase SPCL’s ability to maintain anatomical fidelity in clinically challenging areas, preserving critical vessel connectivity while suppressing background noise, ultimately yielding segmentations that more closely match groundtruth annotations.
Complementing our SPCL evaluation, the second perspective evaluates the proposed MSA-UNet3+ against fifteen state-of-the-art medical segmentation models: SwinUNet [54], CMU_Net [55], MMDC_Net [56], FR_Unet [57], Isunetv1 [46], CA_Net [58], MBS_Net [59], CMU_NeXt [60], BCU_Net [61], MCDAU_Net [45], MGA_Net [62], IMFF_Net [63], DATrans_Unet [64], and PMFS_Net [65]. Through comprehensive quantitative metrics and qualitative visual comparisons, we demonstrate MSA-UNet3+’s superior segmentation performance, particularly in challenging anatomical regions.
Figure 8: Comparison of Deep Learning Models for coronary DSA Image Segmentation: Model Efficiency vs. Segmentation Accuracy. The figure illustrates the trade-off between model efficiency (parameters in millions) and segmentation accuracy ( average DSC, and ASD) for various deep learning models, including the proposed MSA-UNet3+. The size and color of each scatter point represents the model size and ASD values respectively, with the colorbar providing a visual mapping.
Figure 9: Boxplot of Dice coefficient is shown for all test samples of the coronary DSA dataset. In the boxes, the orange line within each box represents the median, and the green triangle represents the mean.

### v3

@@ 3.3.1 Multi-scale attention encoder (M-Encoder)
3.3.2 Multi-scale dilated bottleneck (MSD-Bottleneck)
Continuous convolution and pooling operations frequently degrade fine-grained details crucial for segmentation, while conventional up-sampling fails to recover them. We address this with our MSD-Bottleneck (Fig. 5), building upon [44], combining a bottleneck structure with an Atrous Spatial Pyramid Pooling (ASPP) to preserve fine details while capturing multi-scale context. It uses three dilation patterns ([1,2,1], [2,4,2], [4,8,4]) to extract multi-scale features efficiently. Residual connections prevent vanishing gradients, while ASPP’s parallel dilated convolutions (rates=[4,8]) and global average pooling enhance context. Features are concatenated and projected to the original channel dimension, combining local precision with global context. For the input $x$, processing follows Eq. (2): first through dilated bottleneck blocks with residuals, then ASPP.
@@ 3.3.2 Multi-scale dilated bottleneck (MSD-Bottleneck)
$\text{out}_{2}=\text{conv}_{3\times 3}(\text{out}_{1};\text{dilation}=[2,4,2]),$
$\text{out}_{3}=\text{conv}_{3\times 3}(\text{out}_{2};\text{dilation}=[4,8,4]),$
$\text{out}_{4}=\text{Cat}\Big{(}(\text{out}_{3};\text{dilation}=[4,8]),\text{conv}_{1\times 1}(\text{AAP}(\text{out}_{3}))\Big{)},$
@@ 3.3.3 Contextual attention fusion module (CAFM)
Encoders in segmentation networks lose spatial context in deeper layers, particularly in complex medical imaging such as coronary digital subtraction angiography (DSA). Our Contextual Attention Fusion Module (CAFM) (Fig. 6) addresses this by multi-scale dilated convolutions (rates=[1,2,4,8]) to expand receptive fields while preserving resolution, enabling robust feature extraction across varying vessel sizes; a key requirement in coronary DSA. Furthermore, CAFM incorporates Squeeze-and-Excitation (SE) blocks to dynamically recalibrate channel-wise feature importance, prioritizing diagnostic features. CAFM provides the decoder with refined, context-rich inputs, bridging encoder-decoder semantic gaps through enhanced multi-scale representations, this improves segmentation of fine and coarse structures. Let $x$ be the input feature map to the CAFM module. The operations performed by the module can be described as follows:
$\text{out}_{\text{final}}=\text{conv}_{3\times 3}\Big{(}SE\big{(}Cat(\text{x};\text{dilation}=[1,2,4,8])\big{)}\Big{)}$ (3)
@@
Figure 6: The architecture of the Contextual Attention Fusion Module (CAFM) in the MSA-UNet3+ model, combines Squeeze-and-Excitation (SE) blocks with multi-scale processing.
@@ 3.3.4 Loss function
In this study, we optimize our deep learning model for coronary DSA image segmentation using a composite loss function addressing class imbalance and intra-class variability. While binary cross-entropy (BCE) and Dice loss are commonly employed, they do not inherently enforce semantically meaningful or discriminative feature learning in the encoder. To address this limitation, we combine them with the proposed novel supervised prototypical contrastive loss (SPCL), which encourages the encoder to generate semantically rich and discriminative representations, thereby enhancing the overall performance of the segmentation network.
@@
    It measures the dissimilarity between the predicted probability map and the ground truth labels. BCE loss ensures the model learns to classify each pixel correctly by penalizing deviations from the ground truth. However, it may struggle with class imbalance, as it treats all pixels equally, regardless of class distribution. For a binary segmentation task, the BCE loss is defined as in Eq.  (4):
@@
    where $y_{i}$ is the ground truth label (0 or 1) for pixel $i$, $\hat{y}_{i}$ is the predicted probability of pixel $i$ belonging to the foreground class, and $N$ is the total number of pixels.
@@
    Dice loss quantifies the similarity between the predicted and ground truth segmentation masks. By maximizing this overlap, it directs the model’s focus to regions of interest, enhancing segmentation performance. Dice loss complements BCE by prioritizing accurate foreground class predictions, essential for precise segmentation in coronary DSA images. The Dice loss is defined in Eq.  (5):
@@
    where $y_{i}$ and $\hat{y}_{i}$ are the ground truth and predicted probability for pixel $i$, respectively.
@@
    To address class imbalance and high intra-class variance, we introduce the Supervised Prototypical Contrastive Loss (SPCL). This loss combines supervised and prototypical contrastive losses, enhancing the encoder’s ability to differentiate between diverse classes. (1) Supervised Contrastive Loss: This loss encourages the encoder to map feature vectors of similar image samples (foreground or background regions) closer in the embedding space while pushing dissimilar ones apart. The loss function is defined as in Eq. (6):
@@
    where $f_{i}$ and $f_{j}$ are feature vectors, $P(i)$ is the set of positive pairs for $i$, $A(i)$ is the set of all pairs for $i$, sim is a cosine similarity function, and $\tau$ is a temperature parameter set to 1 without hyperparameter tuning.
@@
    where $N$: Total number of valid embeddings (pixels), $n_{\text{p}}$: Number of prototypes set to 2 without hyperparameter tuning, $D\,(z_{i},p_{k})$: Cosine distance between embedding $z_{i}$ and prototype $p_{k}$, $w_{1}$: Weight for positive samples (close to their prototype), $w_{0}$: Weight for negative samples (far from their prototype), and $m$: Margin parameter controlling class separation.
@@ 4.1 Dataset
The DSA dataset used in this study was acquired from the authors of [45], with original data provided by the General Hospital of Southern Theater Command. The dataset consists of 300 coronary X-ray angiographic sequences from 50 patients, each with a resolution of $512\times 512$ pixels, including both left and right coronary imaging sessions. For this study, we utilized only the right coronary artery images (150 images), which capture contrast agent flow through the coronary vasculature. Clinicians selected the most diagnostically relevant frames, which were then resized to 256$\times$256 pixels during pre-processing. The dataset was divided into 120 training and 30 testing images. We employed 5-fold cross-validation on the training set, with results averaged across all folds for final evaluation.
@@ 4.2 Implementation details
The proposed model was implemented using the PyTorch and PyTorch Lightning frameworks, enabling flexible and reproducible deep learning experimentation. All experiments were conducted on an NVIDIA RTX 4090 GPU (24GB VRAM). Each image and its corresponding label were resized to a fixed resolution of 256×256 pixels. The dataset, comprising $256\times 256$ pixel images and corresponding labels, was normalized to the range $[-1,1]$ using a mean and standard deviation of 0.5 [46]. During training, we applied data augmentation techniques, including random brightness [47] and contrast adjustments [48], to improve model generalization. Optimization was performed using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$, decayed by a factor of 0.1 at 60% and 80% of the total 100 training epochs. A batch size of 5 was maintained throughout training.
@@ 4.3 Evaluation metrics
    It measures the overlap between the predicted segmentation mask and the ground truth, it emphasizes correct foreground class predictions. The mathematical representation of DSC is illustrated in Eq. ( 9):
@@
    where True Positives (TP): Pixels correctly predicted as foreground; False Positives (FP): Pixels incorrectly predicted as foreground; and False Negatives (FN): Pixels incorrectly predicted as background.
@@
  3. 3.
    F1-score
[…]
    Precision measures the proportion of correctly predicted foreground pixels ($\text{Precision}=\frac{TP}{FP+TP}$). The F1-score balances precision and recall, making it a comprehensive metric particularly in imbalanced datasets.
  4. 4.
    Average surface distance (ASD)
    ASD measures the average distance between the surfaces of the predicted and ground truth masks, as defined in Eq.  (12):
    where, $S_{P}$ and $S_{G}$ are the surfaces of the predicted and ground truth masks, respectively; $d(p,S_{G})$ is the minimum Euclidean distance from point $p$ on $S_{P}$ to $S_{G}$; and $d(g,S_{P})$ is the minimum Euclidean distance from point $g$ on $S_{G}$ to $S_{P}$.
    ASD evaluates segmentation boundary accuracy, with lower values indicating better alignment between predicted and ground truth contours, crucial for tasks such as coronary artery segmentation.
  5. 5.
    Average contour distance (ACD)
    ACD measures the average distance between the contours of the predicted and ground truth masks, specifically focusing on contour points. It is defined as in Eq. (13):
Table 1: Segmentation performance comparison on coronary DSA dataset assessing SPCL effectiveness. Metrics shown for six architectures at $\gamma=0$ and $\gamma=1$, with 5-fold cross-validation results (means $\pm$ standard deviations).
@@
Figure 8: Visualization comparison of R2AU-Net (blue border) and R2U-Net (orange border) segmentation with/without SPCL loss on coronary DSA images. Left to right: original angiogram, ground truth,$\gamma=0$, and SPCL-enhanced results ($\gamma=1$). Yellow rectangles highlight SPCL’s improvements in vessel continuity and false positive reduction.
@@ 4.4 Comparison results
As shown in Table 1, our 5-fold cross-validated results on the coronary DSA dataset demonstrate consistent and statistically significant improvements when employing SPCL. Specifically, the Dice coefficient increases from 87.33%, 87.49%, 86.81%, 86.21%, 87.15%, and 85.68% to 87.87%, 87.85%, 87.83%, 87.03%, 87.79%, and 86.80% for U-Net, Att U-Net, UNet++, R2U-Net, Attention UNet++, and R2AU-Net respectively. These quantitative improvements across all architectures and metrics confirm SPCL’s effectiveness as a generalizable enhancement for medical image segmentation tasks. Our comparative analysis of loss functions for coronary DSA segmentation (Fig. 7) reveals that integrating SPCL with Dice+BCE yields consistent improvements in all architectures, with DSC gains ranging from +0.36 (U-Net) to +1.12 (R2AU-Net). This performanance progression aligns with the findings presented previouusly in Table 1, confirming the dual ability of SPCL to: (1) improve feature discrimination through semantic clustering (2) address class imbalance through contrastive learning. The demonstrated robustness across architectures (U-Net to R2AU-Net) and the magnitude of improvement position SPCL as a general solution for medical segmentation tasks.
Table 2: Performance comparison of state-of-the-art medical image segmentation models on the coronary DSA dataset. Results are presented as (means $\pm$ standard deviations) across 5-fold cross-validation. Statistical significance of performance improvements achieved by MSA-UNet3+ is verified using a one-tailed $t$-test.
@@
Figure 9: Comparison of Deep Learning Models for Coronary DSA Image Segmentation: Model Efficiency vs. Segmentation Accuracy. The figure illustrates the trade-off between model efficiency (parameters in millions) and segmentation accuracy (average DSC and ASD) for various deep learning models, including the proposed MSA-UNet3+. The size and color of each scatter point represent the model size and ASD values, respectively, with the color bar providing a visual mapping.
Figure 10: A Boxplot of the Dice coefficient is shown for all test samples of the coronary DSA dataset. In the boxes, orange line and green triangle within each box represents the median and the mean, respectively.
These results corroborate our earlier quantitative assessments while providing new insights into SPCL’s architecture-dependent optimization characteristics. The qualitative comparison in Fig. 8 demonstrates the effectiveness of SPCL through two key improvements: (1) enhanced vessel continuity, particularly in complex branching structures where the baseline model fragments vessels, and (2) superior preservation of thin vessels that conventional approaches often miss. These visual improvements align with our quantitative metrics, as illustrated in Table 1. The yellow-highlighted regions specifically showcase SPCL’s ability to maintain anatomical fidelity in clinically challenging areas, preserving critical vessel connectivity while suppressing background noise, ultimately yielding segmentation that more closely matches ground truth annotations.
Complementing our SPCL evaluation, the second perspective evaluates the proposed MSA-UNet3+ with SPCL loss against fifteen state-of-the-art medical segmentation models: SwinUNet [53], CMU_Net [54], MMDC_Net [55], FR_Unet [56], Isunetv1 [45], CA_Net [57], MBS_Net [58], CMU_NeXt [59], BCU_Net [60], MCDAU_Net [44], MGA_Net [61], IMFF_Net [62], DATrans_Unet [63], and PMFS_Net [64]. Through comprehensive quantitative metrics and qualitative visual comparisons, we demonstrate MSA-UNet3+’s superior segmentation performance, particularly in challenging anatomical regions.